OpenStudy (anonymous):

Does anyone know how much kg is 800 grams?

OpenStudy (anonymous):

0.800 kg!

OpenStudy (anonymous):

kilo is prefix denoting 1000. a kilogram is 1000 grams. so 800 grams is 0.8 kg

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

how much yards in 18 feet

OpenStudy (anonymous):

?????????

OpenStudy (accessdenied):

How many feet are in one yard?

OpenStudy (anonymous):

1 yard = 3 feet

OpenStudy (accessdenied):

We know that \( 1 \; \text{yard} = 3 \; \text{feet} \). This makes the following true: \( \displaystyle \frac{1 \; \text{yard}}{3 \; \text{feet}} = 1 \), \( \displaystyle \frac{3 \; \text{feet}}{1 \; \text{yard}} = 1 \) Just by dividing both sides. So, 1 is special because we can multiply anything by 1 and have the same thing. Well, if we have: \( 18 \; \text{feet} \) We can multiply by a unit that cancels feet and leaves yards. \( \displaystyle 18 \; \text{feet} \times 1 \) \( \displaystyle 18 \; \text{feet} \times \frac{1 \; \text{yard}}{3 \; \text{feet}} \) \( \displaystyle 18 \; \cancel{\text{feet}} \times \frac{1 \; \text{yard}}{3 \; \cancel{\text{feet}}} \) \( \displaystyle 18 \times \frac{1}{3} \; \text{yards} \) So, we can see that we simply take 18/3 to get our number of yards. This same method will work with basically any unit conversions. :)

OpenStudy (anonymous):

ya I got it

OpenStudy (anonymous):

Approximately how many square centimeters are there in 22 square inches? Round to the nearest tenth

OpenStudy (anonymous):

HELLOO

OpenStudy (anonymous):

IS ANYONE GONNA TUTOR ME OR SOMETHING!!!

OpenStudy (anonymous):

YOU PEOPLE DONT HELP AT ALL DO YOU?

OpenStudy (anonymous):

IM GONNA FIND A DIFFRENT WEBSITE THAT WILL ACTUALLY HELP ME!!!

OpenStudy (marissalovescats):

http://www.unitconverters.net/ This site might help a bit better :)